Update for the Week Ending February 12, 2016

Site Mobilization Completed
Site Remediation Begins

On-going activities:

  • Continued the installation of Site electrical service for the water treatment plant (WTP), TCB, and TCB vapor management system (VMS);
  • Continued the application of water and calcium chloride for dust control;
  • Continued the application of odor-suppressant foam as needed for dust and odor control;
  • Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site);
  • Continued the removal of hardscape (asphalt and concrete) over the Holder No. 2 tank;
  • Continued the excavation of shallow soils within the Holder No. 2 tank beneath the TCB in preparation for the bucket soil mixing (BSM) demonstration test;
  • Initiated operation of the WTP;
  • Initiated truck lining, loading, and shipment of impacted soils to Bayshore Soil Management for treatment and disposal:
  • 12 trucks: 322.06 tons shipped last week; and
  • 322.06 tons shipped to date.
  • Initiated the BSM demonstration test.

Recently completed activities:

  • Completed mobilization of equipment and supplies;
  • Completed installation of the VMS components; and
  • Completed start-up and testing of the WTP.

Upcoming activities:

  • Complete the BSM demonstration test in the Holder No. 2 tank within the TCB;
  • Conduct full scale BSM in the Holder No. 2 tank;
  • Continue the shipment of impacted shallow soils for treatment and disposal;
  • Ship removed asphalt for recycling;
  • Perform pre-trenching for on-Site utility clearance along the support of excavation (SOE) wall perimeter;
  • Decommission former monitoring wells on Site;
  • Install electrical service from Con Edison and demobilize the temporary generator from the Site; and
  • Install dewatering wells, initiate ground water extraction, and begin operation of the WTP.

Community Air Monitoring Plan Summary

Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.

There were no fifteen-minute Action Limit exceedances of Total Volatile Organic (TVOCs) or odors during this reporting period.

There were six instances of the fifteen-minute average concentration of Particulate Matter (PM-10) [dust] on February 11, 2016, exceeding the applicable Action Limit during windy conditions.  These concentrations are attributed to clean fill.  Work was stopped repeatedly and water and calcium chloride were applied until dust levels returned to below the response limit and Site work resumed.

If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.

Photos of Activity

Continuing the excavation of shallow soils in the Holder No. 2 tank within the temporary containment building (TCB)

Continuing the excavation of shallow soils in the Holder No. 2 tank in the TCB

Continuing the transfer of reusable soil to the on-Site stockpile

Excavating material from the Holder No. 2 tank in preparation for the Bucket Soil Mix (BSM) demonstration test

Initiating the BSM demonstration test in the Holder No. 2 tank within the TCB

Loading the lined trucks with impacted soil for shipment to Bayshore Soil Management for treatment and disposal

Cleaning trucks on the decontamination pad prior to leaving the Site on N. 11th Avenue